OpenStudy (anonymous):

Terry can paint a house in 45 hours. His father can paint the same house in 50 hours and his brother can paint it in 40 hours. How long will it take all three working together to do the painting? Help please? I need the solution. THANKS!

OpenStudy (amistre64):

How much can they each do in 1 hour alone?

OpenStudy (amistre64):

Terry takes 45 hours to do the whole thing; so 45/45 is complete for him. How much is done in 1 hour? 1/45 right?

OpenStudy (amistre64):

Dad can do it all in 50 hours: 50/50 is complete for him. How much is done in 1 hour?

OpenStudy (anonymous):

And the brother can do it in 40/40.

OpenStudy (amistre64):

good; so each one alone can do: 1/45 + 1/50 + 1/40, we have to add all those fractions together to be able to solve this now

OpenStudy (anonymous):

then..

OpenStudy (anonymous):

1800/141

OpenStudy (anonymous):

Can u see how ?

OpenStudy (amistre64):

lol .... yeah; I get 121/1800 after adding. then flip it over like rosh has done to get the answer

OpenStudy (anonymous):

no. :'(

OpenStudy (anonymous):

okay. so the answer is 1800/141?

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

In one hour terry can paint 1/45 house Similarly father and brother in 1/50 and 1/40 house respectively So in one hour all can paint 1/45 + 1/50 + 1/40

OpenStudy (anonymous):

how can i convert it in time? i cant answer 1800/121. the answer should be in hours right?

OpenStudy (anonymous):

which will give you 141/1800, now just flip it over to get the answer

OpenStudy (anonymous):

12.76 hours

OpenStudy (anonymous):

sorry should have been 1800/121

OpenStudy (anonymous):

I screwed up on adding LOL, 14.88 hours

OpenStudy (amistre64):

first we get common denominators: I think 5 is a good common denominator ... maybe :) 1/9 1/10 1/8 1/9 + 1/10 + 1/8 ---- + ------ + ----- = ---------------- 45/9 50/10 40/8 5 Now that the numbers are smaller :) we can find a common denominator of 9,10, and 8; 360 pops out in the end. 1*40 1*36 1*45 40+36+45 121 ---- + ----- + ----- = ---------- = ------ 9*40 10*36 8*45 360 360 Now that that is over; lets use it 121/360 121 121 -------- = ------ = ------ ; yay!! 5 5*360 1800

OpenStudy (anonymous):

THANK YOU! GOT IT ALREADY. SO HELPFUL!

OpenStudy (amistre64):

Since there is 1800 "parts" to paint; and 121 get painted in an our we get: 121*hours = 1800 hours = 1800/121 ; which is just the flipped version of what we got prior

OpenStudy (anonymous):

Well done amistre64 Ha Ha

OpenStudy (amistre64):

;) thanx